find the no which when divided by 17 ,16,15 and the remainder is 3.10,0 respectively means u have to find that no which is divided by 17 then it leaves the remainder 3 when divided by 16 leaves remainder 10 and when divi ded by 15 leaves remainder 0


 find the  no which when divided by 17 ,16,15 and the remainder is 3.10,0 respectively  means u have to find


that no which is divided by 17 then it leaves the remainder


3 when divided by 16 leaves remainder 10 and when divi 


ded by 15 leaves remainder 0 


Grade:12

2 Answers

Nikhilesh Reddy
12 years ago

105

Ashwin Muralidharan IIT Madras
12 years ago

Hi Keshav,

 

The answer would be 1210....

 

Method:

N = 17X + 3....

N = 16X + X + 3....

 

Now X+3 = 16Y+10

Also N = 15X + 2X+3

So 2x+3 = 15Z.....

 

Relation between Y and Z is 32Y+17 = 15Z...

So minimum value of Y,Z is 4,9

 

Which would give X = 71....

 

So the smallest number "N" satisfying all the conditions is N=17X+3

=17(71) + 3

=1210

 

Hence 1210 is the answer.

 

Best Regards,

Ashwin (IIT Madras).